OpenStudy (anonymous):

1) Compute the probability that an observation from a N(3, 5) population is less than 7. 2) Find a number Z so that P{ N(4, 4) < Z } = .75 ( the probability that a normal 4, 4 is less than Z equals .75) 3) Compute the probability that a N(14,3) observations is larger than 12. 4) If the heights of 17-your-old girls are known to be normally distributed with mean 64 inches and standard deviation 3 inches, how tall must Taylor be when she is 17 so that she will be taller than 73% of other girls her age?

OpenStudy (paxpolaris):

using Z-table \[\text{1) } \large P(X < 7) = P\left(Z<{7-3\over 5}\right)=P(Z<0.8)\]

OpenStudy (anonymous):

so i look at the z table for 0.8 and that will be the solution

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

ok that makes sense thanks do you happen to know how to do the other ones

OpenStudy (anonymous):

for number 4 i did it and got this numbers but i dont remember how i did it or if they are right they are p(x<65.8384)=P(z,0.6128=0.73

OpenStudy (paxpolaris):

4) 65.84in is right

OpenStudy (anonymous):

would it be hard to try to show me the steps cause to tell you the truth i cant remember how i got there

OpenStudy (paxpolaris):

first find the z value that gives probability 0.73...

OpenStudy (paxpolaris):

how did you get 0.6128?

OpenStudy (anonymous):

um im not sure exactly

OpenStudy (paxpolaris):

i had guessed 0.612 from table, then: \[\Large {X-64 \over 3 }=0.6128\]

OpenStudy (anonymous):

but is it right cause i dont see 0.6128 in the table

OpenStudy (paxpolaris):

you know its between 0.61 and 0.62 ... which is good enough if you need your answer upto 2 decimal places

OpenStudy (paxpolaris):

also you could use invcdf / invnorm function on ti-calculators if you are allowed

OpenStudy (anonymous):

i dont have one those calculators i just havve the basic ones but ya i rememred now were o.6128 came thanks for reminding me

OpenStudy (anonymous):

im still not getting for some reason 65.8384 for my answer

OpenStudy (paxpolaris):

0.6128*3 +64

OpenStudy (anonymous):

oh i was adding 64+0.6128/3 instead of multiplying

OpenStudy (paxpolaris):

for the rest exact answers are: 2)6.697959001 3)0.747507462

OpenStudy (anonymous):

woul it be hard for you to show me the steps cause i have another promblem like this that i have to do and i want to learn the stuff if not its fine

OpenStudy (paxpolaris):

2)same as 4) find the z-value for P=0.75 then *4 + 4

OpenStudy (paxpolaris):

3)\[\Large = 1-P \left( z<{12-14 \over 3} \right)\]

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

wait wouldnt it be 12+14 since its larger than

OpenStudy (paxpolaris):

\[\Large P(p<X)=P \left( z<{X-\mu \over \sigma} \right)\]

OpenStudy (paxpolaris):

12 < 14 so, z< 0 ... the mean of the standard normal function

OpenStudy (anonymous):

ok that makes sense thanks so much your the best

OpenStudy (paxpolaris):

no prblm :)